16=10/3a

15/7y=3

14=-7/9b

can you please show me so i can help my son i'm so rusty with algebra



Answer by stanbon(48535)

1st problem:
Multiply both sides of the equation by 3a to get the following:
48a = 10
Divide both sides by 48 to get:
a = 10/48 = 4/24
2nd problem
Multiply both sides by 7y to get:
15 = 21y
Divide both sides by 21 to get
y = 15/21 = 5/7
3rd problem:
14 = -7/9b
Multiply both sides by 9b to get:
9b(14)= -7
Divide both sides by 7 to get:
9b(2) = -1
18b = -1
Divide both sides by 18 to get:
b = -1/18
Good luck with the refresher.
Cheers,
Stan H.